OpenStudy (anonymous):

i forgot what ! means so i dont remember hot to solve this... Simplify 5!/3!

OpenStudy (anonymous):

\[n!=\prod_{1 \leq f \leq n}f=1\cdot 2\cdot 3\cdots n\]

OpenStudy (vishweshshrimali5):

! means factorial n! = n(n-1)(n-2).........1 now you can solve for 5!/3! = 5*4*3*2*1/3*2*1 = 5*4 = 20

Parth (parthkohli):

\( \color{Black}{\Rightarrow n! = n \times (n -1) \times (n - 2) \times .....\times 1 }\)

OpenStudy (maheshmeghwal9):